# DocParseNet: Advanced Semantic Segmentation and OCR Embeddings for Efficient Scanned Document Annotation

### v2

@@ Abstract
Automating the annotation of scanned documents is challenging, requiring a balance between computational efficiency and accuracy. DocParseNet addresses this by combining deep learning and multi-modal learning to process both text and visual data. This model goes beyond traditional OCR and semantic segmentation, capturing the interplay between text and images to preserve contextual nuances in complex document structures. Our evaluations show that DocParseNet significantly outperforms conventional models, achieving mIoU scores of 49.12 on validation and 49.78 on the test set. This reflects a 58% accuracy improvement over state-of-the-art baseline models and an 18% gain compared to the UNext baseline. Remarkably, DocParseNet achieves these results with only 2.8 million parameters, reducing model size by approximately 25 times and speeding up training by 5 times compared to other models. These metrics, coupled with a computational efficiency of 0.034 TFLOPs (BS=1), highlight DocParseNet’s high performance in document annotation. The model’s adaptability and scalability make it well-suited for real-world corporate document processing applications. The code is available at https://github.com/ahmad-shirazi/DocParseNet.
@@
Keywords: Semantic Segmentation, Multi-Modal Learning, Document Annotation, Optical Character Recognition (OCR), Contextual Understanding, Document Processing, Model Efficiency
@@ 1 Introduction
This gap is further widened by the complex nature of documents, where the structure of characters and their contextual meaning are intricately intertwined. Conventional methods like document layout analysis offer structural parsing (Shen et al., 2021), while semantic analysis aims to uncover deeper meanings (Mercha & Benbrahim, 2023). Although template-based extraction efficiently harvests data from predefined fields in structured documents, these methods are insufficient for documents blending structured and unstructured data. The variability in document quality, format diversity, and layout complexity further challenges current methods, especially with documents exhibiting high variance in text density, font styles, and graphical elements, as well as poor scanning quality (Zhou et al., 2023). Additionally, the computational demand and need for large, annotated datasets for training models exacerbate these challenges, creating a bottleneck for efficient data extraction and processing (Lepcha et al., 2023; Liang et al., 2022).
@@ 2 Related Work
The UNet framework employs an encoder-decoder architecture with skip connections to preserve spatial information (Huang et al., 2020). This approach benefits from convolutional neural networks’ (CNNs’) local feature extraction and transformers’ global perspective. Inspired by the UNet framework, UNeXt (Valanarasu & Patel, 2022) has provided a less computationally intensive model by utilizing a Multi-Layer perception (MLP) approach (Taud & Mas, 2018).
@@ 3 DocParseNet
Our proposed model, DocParseNet (Figure 1), is a multi-modal deep learning architecture designed for parsing and annotating corporate agreement PDFs. It processes images derived from these PDFs, identifying key elements such as ”State,” ”County,” ”Agreement Title,” ”Grantee Company,” and ”Grantor Company.” We trained DocParseNet on a dataset of 4,102 expert-annotated corporate agreement PDFs, comprising 19,821 images, split into training, validation, and testing sets using an 8-1-1 ratio. The non-static positioning of fields, influenced by the dynamic document structure, adds further complexity to the model’s task. This heterogeneity, however, is essential, enriching the dataset and providing a robust foundation for efficient DocParseNet training.
@@ 3.1 DocParseNet_UNet
CNNConv 2D+Mish: The CNN component of DocParseNet_UNet as demonstrated by equation 1 uses a series of 2D convolutional layers to extract hierarchical spatial features from input images, utilizing a 1x1 convolution (conv1x1) to refine feature maps between stages. The CNN employs the Mish activation function, which preserves more information compared to traditional activation functions like ReLU. Equation 2 takes the input image (I) and outputs a random crop (I). The random crop is then passed onto the aforementioned CNN component, and the output is calculated.
@@
Shifted MLP Encoding Block: The Shifted MLP Encoding Block is designed to overcome the limitations of CNNs, particularly their localized receptive fields. By employing a shift operation in the MLP’s receptive window across the feature map, it achieves cross-window connectivity, capturing broader contextual information efficiently. This structure enriches feature representation with contextual information spanning both local and distant regions of the input. To briefly define the shifted MLP, we have:
$V_{n}=\text{Linear}\left(\text{Mish}\left(\text{DWConv}\left(W_{1}\cdot X_{\text{shift}}+b_{1}\right)\right)\right)$ (3)
Equation 3 summarizes the shifted MLP, where a Mish activation function is applied after Depth-Wise Convolution (DWConv) and the feed-forward neural network. Another feed-forward neural network is then applied to the output. Additional details are provided in Appendix A.1.
@@ 3.2 DocParseNet_OCR
The OCR phase in DocParseNet, shown in Figure 1, is designed for text recognition within document images. It uses optical character recognition to convert visual text into machine-encoded text, enabling the extraction of textual embeddings. This study utilizes a pretrained DistilBERT model for its efficient text encoding capabilities. The OCR process includes Image-to-Text Conversion ($f_{\text{OCR}}$), text embedding using the Distilbert model ($f_{\text{BERT}}$), and [CLS] Token Embedding (E). For further details, please refer to Appendix A.2.
@@ 3.3 DocParseNet_Fusion
In DocParseNet, the fusion process combines the shifted MLP encoding block output with BERT embeddings via multi-head attention:
- Let $V_{6}\in\mathbb{R}^{B\times C_{6}\times\frac{H}{32}\times\frac{W}{32}}$ be the output from the shifted MLP encoding block.
@@
The multi-head attention outputs a feature map $F\in\mathbb{R}^{B\times C^{\prime}_{6}\times\frac{H^{\prime}}{32}\times\frac{W^{\prime}}{32}}$.
$F=f_{\text{MultiHead}}(V_{6},E)$ (5)
@@ 4 Result
In our empirical study, we evaluated the proposed DocParseNet model by benchmarking it against leading semantic segmentation models, including GCNet (Cao et al., 2020), UPerNet (Xiao et al., 2018), DeeplabV3 (Chen et al., 2017), and Segformer (Xie et al., 2021), using the segmentation (Contributors, 2020) framework. These state-of-the-art models, optimized for image feature extraction, do not inherently account for text-based information.
The comparative analysis, detailed in Table 1, Appendix B.2 (Table B1 and Figure B2), highlights DocParseNet’s enhanced ability to incorporate text information, resulting in a significant performance boost. Our model’s superior performance is evidenced by its mean Intersection over Union (mIoU) scores: 49.12 on the validation set and 49.78 on the test set. This demonstrates its robustness and superior generalization in handling the challenges of semantic segmentation and OCR embedding for scanned document annotation. Additionally, the training dynamics of DocParseNet, as shown in Appendix B.1, Figure B1, exhibit robust learning and generalization capabilities through a consistent upward trajectory in IoU scores and a decrease in loss, highlighting its resistance to overfitting.

### v3

@@ Abstract
Automating the annotation of scanned documents is challenging, requiring a balance between computational efficiency and accuracy. DocParseNet addresses this by combining deep learning and multi-modal learning to process both text and visual data. This model goes beyond traditional OCR and semantic segmentation, capturing the interplay between text and images to preserve contextual nuances in complex document structures. Our evaluations show that DocParseNet significantly outperforms conventional models, achieving mIoU scores of 49.12 on validation and 49.78 on the test set. This reflects a 58% accuracy improvement over state-of-the-art baseline models and an 18% gain compared to the UNext baseline. Remarkably, DocParseNet achieves these results with only 2.8 million parameters, reducing model size by approximately 25 times and speeding up training by 5 times compared to other models. These metrics, coupled with a computational efficiency of 0.039 TFLOPs (BS=1), highlight DocParseNet’s high performance in document annotation. The model’s adaptability and scalability make it well-suited for real-world corporate document processing applications. The code is available at https://github.com/ahmad-shirazi/DocParseNet.
@@
Keywords: Document AI, Multi-Modal Learning, Vision-and-Language, Document Annotation, document image understanding
@@ 1 Introduction
This gap is further widened by the complex nature of documents, where the structure of characters and their contextual meaning are intricately intertwined. Conventional methods like document layout analysis offer structural parsing (Shen et al., 2021; Zhou & Ramnath, 2022), while semantic analysis aims to uncover deeper meanings (Mercha & Benbrahim, 2023). Although template-based extraction efficiently harvests data from predefined fields in structured documents, these methods are insufficient for documents blending structured and unstructured data. The variability in document quality, format diversity, and layout complexity further challenges current methods, especially with documents exhibiting high variance in text density, font styles, and graphical elements, as well as poor scanning quality (Zhou et al., 2023). Additionally, the computational demand and need for large, annotated datasets for training models exacerbate these challenges, creating a bottleneck for efficient data extraction and processing (Lepcha et al., 2023; Liang et al., 2022).
@@ 2 Related Work
The UNet framework employs an encoder-decoder architecture with skip connections to preserve spatial information (Huang et al., 2020; Navard & Yilmaz, 2024). This approach benefits from convolutional neural networks’ (CNNs’) local feature extraction capabilities. Inspired by the UNet framework, UNeXt (Valanarasu & Patel, 2022) offers a less computationally intensive model by utilizing a Multi-Layer Perceptron (MLP) approach (Taud & Mas, 2018).
@@ 3 DocParseNet
Our proposed model, DocParseNet (Figure 1), is a multi-modal deep learning architecture designed for parsing and annotating corporate agreement PDFs. It processes images derived from these PDFs, identifying key elements such as State, County, Agreement Title, Grantee Company, and Grantor Company. We trained DocParseNet on a dataset of 4,102 expert-annotated corporate agreement PDFs, comprising 19,821 images, split into training, validation, and testing sets using an 8-1-1 ratio. The non-static positioning of fields, influenced by the dynamic document structure, adds further complexity to the model’s task. This heterogeneity, however, is essential, enriching the dataset and providing a robust foundation for efficient DocParseNet training.
@@ 3.1 DocParseNet_UNet
CNNConv 2D+Mish: The CNN component of DocParseNet_UNet as demonstrated by equation 1 uses a series of 2D convolutional layers to extract hierarchical spatial features from input images, utilizing a 1x1 convolution (conv1x1) to refine feature maps between stages. The CNN employs the Mish activation function (Misra, 2019), which preserves more information compared to traditional activation functions like ReLU. Equation 2 takes the input image (I) and outputs a random crop (I). The random crop is then passed onto the aforementioned CNN component, and the output is calculated.
@@
Shifted MLP Encoding Block: The Shifted MLP Encoding Block is designed to overcome the limitations of CNNs, particularly their localized receptive fields. By employing a shift operation in the MLP’s receptive window across the feature map, it achieves cross-window connectivity, capturing broader contextual information efficiently. This structure enriches feature representation with contextual information spanning both local and distant regions of the input. The shifted MLP can be described as the following:
$V_{n}=\text{Linear}\left(\text{Mish}\left(\text{DWConv}\left(W_{1}\cdot X_{\text{shift}}+b_{1}\right)\right)\right),$ (3)
where a Mish activation function is applied after Depth-Wise Convolution (DWConv) and a feed-forward neural network. Another feed-forward neural network is then applied to the output. Additional details are provided in Appendix A.1.
@@ 3.2 DocParseNet_OCR
The OCR phase in DocParseNet, shown in Figure 1, is designed for text recognition within document images. It uses optical character recognition to convert visual text into machine-encoded text, enabling the extraction of textual embeddings. This study utilizes a pre-trained DistilBERT model for its efficient text encoding capabilities. The OCR process includes Image-to-Text Conversion ($f_{\text{OCR}}$), text embedding using the Distilbert model ($f_{\text{BERT}}$), and [CLS] Token Embedding (E). For further details, please refer to Appendix A.2.
@@ 3.3 DocParseNet_Fusion
In DocParseNet, the fusion process combines the last shifted MLP encoding block output with BERT embeddings via multi-head attention:
- Let $V_{6}\in\mathbb{R}^{B\times C_{6}\times\frac{H}{32}\times\frac{W}{32}}$ be the output from the last shifted MLP encoding block.
@@
The multi-head attention outputs a feature map $F\in\mathbb{R}^{B\times C^{\prime}_{6}\times\frac{H^{\prime}}{32}\times\frac{W^{\prime}}{32}}$ by:
$F=f_{\text{MultiHead}}(V_{6},E).$ (5)
@@ 4 Result
In our empirical study, we evaluated the proposed DocParseNet model by benchmarking it against leading semantic segmentation models, including GCNet (Cao et al., 2020), UPerNet (Xiao et al., 2018), DeeplabV3 (Chen et al., 2017), and Segformer (Xie et al., 2021), using the segmentation (Contributors, 2020; Perera et al., 2024) framework. These state-of-the-art models, optimized for image feature extraction, do not inherently account for text-based information.
The comparative analysis, detailed in Table 1 and Appendix B.2 (Table B1 and Figure B2), highlights DocParseNet’s enhanced ability to incorporate text information, resulting in a significant performance boost. Our model’s superior performance is evidenced by its mean Intersection over Union (mIoU) scores of 49.12 on the validation set and 49.78 on the test set. This demonstrates its robustness and superior generalization in handling the challenges of semantic segmentation and OCR embedding for scanned document annotation. Additionally, the training dynamics of DocParseNet, as shown in Appendix B.1, Figure B1, exhibit robust learning and generalization capabilities through a consistent upward trajectory in IoU scores and a decrease in loss, highlighting its resistance to overfitting.